# Terrier: A Deep Learning Repeat Classifier

### v1

@@ 5.1 Rice
The Rice dataset contains 75 TE models. Using an NVIDIA H100 GPU, Terrier classified this dataset in a total wall time of 6.3s. With a threshold of 0.7, Terrier classified 71 out of the 75 models to the ‘Order’ level, of which 67 were correctly classified, resulting in an accuracy of 94.37%. This accuracy surpassed the best result from the other packages, with TEclass2 achieving 86.9%. Raising Terrier’s classification threshold to 0.9 improved accuracy to 97.1%, but reduced the number of classified sequences to 68. At the ‘Superfamily’ level, Terrier classified over 90% (68 out of 75) of sequences, with only one incorrect classification, resulting in an accuracy of 98.5%. This was much higher than the best result for the other packages, which was 82% for TEclass2.
@@ 5.2 Fruit Fly
The Fruit Fly dataset contains 667 TE models. Terrier classified this dataset in 9.74s, using the same GPU mentioned above. Using a threshold of 0.7, Terrier classified 80.8% of the sequences at the ‘Order’ level, of which 87.9% were correct. This outperformed all other software packages, including TEclass2, which at a threshold of 0.9, achieved an accuracy of 86.4%, but only classified 45.3% of sequences. Using a higher threshold of 0.9 raises the accuracy to 94.5%, but reduces the number of sequences classified to 65.4%. At the ‘Superfamily’ level, using a threshold of 0.7, Terrier predicted 91% of sequences accurately, surpassing TEclass2 at a higher threshold of 0.9, while classifying a far higher proportion of sequences (63.6% compared with 45.2% for TEclass2). When the threshold was raised to 0.9, Terrier’s accuracy increased to 96.6%, but the number of sequences classified decreased to 53.7%.
@@
Figure 9: Comparison between results from RepeatModeler (left) and Terrier (right) experimental data of 8 flatworms and 51 amphibians.
@@ 6 Experimental Data
We present the overall computation time for running Terrier, TERL and DeepTE on these input files. We tested timings on CPUs (eight Intel Xeon Gold 6254) and GPU (NVIDIA A100) (Fig. 8). In all cases, file size was linearly correlated with computational time. Using the CPUs, Terrier scaled at approximately 72s/MB. DeepTE performed faster at around 44s/MB and TERL was faster still at 2.9s/MB. Terrier performed more than 9 times faster with a GPU relative to using the CPUs, scaling at 7.7s/MB. This was faster than DeepTE at 21.3s/MB but still not as fast as TERL which was extremely fast at 0.1s/MB.
In Fig. 9 we show the Order classifications from both RepeatModeler and Terrier at a threshold of 0.7. The RepeatModeler output gives 73.5% as ‘Unknown’ whereas the Terrier classifications are only 34.4% ‘Unknown’.
@@ 7 Conclusion
In this study, we introduce Terrier, a powerful software package designed for the precise classification of repeats from DNA sequences. Terrier distinguishes itself by incorporating a significantly greater number of prediction categories than other comparable methods, providing a finer level of detail for repeat classification. This expanded classification system enables more accurate and comprehensive assessments, especially when dealing with complex, diverse, or poorly characterized genomic regions.
@@
Compared to other state-of-the-art deep learning classifiers, Terrier demonstrates superior performance in terms of both classification accuracy and the breadth of sequences it can classify. The tool’s high precision, combined with its broad applicability across a range of species and repeat types, positions it as a leading tool for repeat classification. These capabilities make Terrier an essential tool for researchers studying repeat-driven evolution, genomic instability, and other areas where a detailed understanding of repeats is crucial. By advancing the field of repeat classification, Terrier enables more accurate genomic annotations and supports the exploration of complex genomic features in both model and non-model organisms.
@@ 8 Data availability
The software is available under an Apache 2.0 Open Source Licence and can be downloaded and installed from GitHub [^7] https://github.com/rbturnbull/terrier. Weights for the model are available with the 0.2.0 release of the software and this is automatically downloaded when first using the model. We also provide a notebook which can be launched in Google Colab to run Terrier on a cloud GPU and to download the results.
The data underlying this article are available in FigShare, at https://doi.org/10.26188/28578992.
@@ 10 Author contributions statement
R.T. conceived and wrote the Terrier software, performed the training. R.T., N.Y. and T.K. prepared the training dataset. R.T. and E.T. ran the software packages on the test datasets. N.Y. prepared the flatworm experimental data. T.K. and E.T. prepared the amphibian experimental data. All reviewed the manuscript.
@@ 11 Acknowledgments
This research was supported by The University of Melbourne’s Research Computing Services and by Australian Research Council grants FT190100462 and LP200301370 awarded to L.F.S. We acknowledge the help of Priyanka Pillai, Swetha Gopikumar Sreeja and Rafsan Al Mamun.

### v2

@@ 5.1 Rice
The rice dataset contains 75 TE models. Terrier classified it in 7.2s. With a threshold of 0.7, 71 models (94.4%) were classified at the ‘Order’ level with 94.4% accuracy, outperforming all other tools. Raising the threshold to 0.9 slightly reduced classified sequences (68) but improved accuracy to 97.1%. At the ‘Superfamily’ level, Terrier classified 68 sequences with 98.5% accuracy—substantially higher than TEclass2’s best of 82.0%. TERL was able to classify more families but the accuracy was substantially.
@@ 5.2 Fruit Fly
This dataset includes 667 TE models, classified in 10.7s. At a 0.7 threshold, Terrier achieved 80.8% classification at the ‘Order’ level with 87.9% accuracy. A threshold of 0.9 raised accuracy to 94.5%, with a drop in classified sequences to 65.4%. At the ‘Superfamily’ level, Terrier reached 91.0% accuracy at 0.7 and 96.6% at 0.9—both outperforming TEclass2 while classifying similar numbers of sequences.
5.3 Human
The human dataset has 1613 TE models, classified in 17.4s. At 0.7, Terrier classified 85.3% of models at the ‘Order’ level with 89.5% accuracy; increasing the threshold to 0.9 raised accuracy to 94.4% (78% classified). At the ‘Superfamily’ level, Terrier showed slightly higher accuracy than TEclass2 at the equivalent thresholds, while classifying more models.
5.4 Mouse
The mouse dataset has 1779 models and was classified in 18.0s. At a threshold of 0.9, it achieved the highest ‘Order’ accuracy (95.6%) while maintaining broad coverage (78.5%). At the ‘Superfamily’ level, Terrier outperformed TEclass2 at the equivalent thresholds, again with broader coverage.
Scatter plots showing the computation time for running Terrier, DeepTE and TERL against the filesize of the input data. The left scatter plot shows timing on just CPUs. The right scatter plots show timings using a GPU. All software packages scale linearly according to filesize. DeepTE and TERL are faster than Terrier using just CPUs but when using a GPU, Terrier is faster than DeepTE whilst still being slower than TERL.
@@ 5.4 Mouse
Bar charts comparing the results from RepeatModeler (left) and Terrier (right) on the experimental data of 8 flatworms and 51 amphibians. The bar charts visualise the number of DNA, LINE, LTR, Other and Unknown classifications from both software packages. The number of Unknown predictions in the Terrier results is less than the corresponding results from RepeatModeler.
Figure 8: Comparison between results from RepeatModeler (left) and Terrier (right) on the experimental data of 8 flatworms and 51 amphibians. The percentage of sequences classified as ‘Unknown’ is labeled for each species.
@@ 6 Experimental Data
We present the overall computation time for running Terrier, TERL and DeepTE on these input files. We tested timings on an GPU (with two CPUs) and then with two CPUs with no GPU (Fig. 7). In all cases, file size was linearly correlated with computational time. Using the CPUs, Terrier scaled at approximately 166s/MB. DeepTE performed faster at around 46s/MB and TERL was faster still at 6.5s/MB. Terrier performed many times faster with a GPU relative to using the CPUs, scaling at 6.8s/MB. This was faster than DeepTE at 21.3s/MB but still not as fast as TERL which was extremely fast at 0.1s/MB. These timings show that Terrier is able to produce classifications for even large genomes within a reasonable timeframe, especially when a GPU is available.
In Fig. 8 we show the Order classifications from both RepeatModeler and Terrier at a threshold of 0.7. The RepeatModeler output gives 73.5% as ‘Unknown’ whereas the Terrier classifications, this number is reduced to only 34.4% ‘Unknown’. The substantial reduction in the number of unclassified repeat families is consistent across amphibians and flatworms of different species and genome sizes.
@@
These experimental results across a wide range of flatworm and amphibian species demonstrate that Terrier can efficiently classify transposable elements in large, repeat-rich genomes. Terrier considerably reduces the number of unclassified repeat families compared to RepeatModeler alone, enabling more comprehensive downstream analysis.
6.1 Northern Krill
To further demonstrate use-cases for Terrier, we applied it classifying unknown TE libraries in northern krill (Meganyctiphanes norvegica)—a large genome of more than 19 Gb. Unneberg et al. Unneberg et al. (2024) recently found that repeats account for 74% of the genome. They released a library of 10909 distinct repeat sequences, of which 1292 (11.8%) were unclassified. We used Terrier to classify these unclassified sequences (Fig. 9). With the default threshold of 0.7, Terrier classified 626 (48.5%) to at least the Order level with 162 to the Superfamily level. These repeat families correspond to more than 10 million individual repeats summing to approximately 2.5 Gb. With the more restrictive threshold of 0.9, Terrier classified 337 (26.1%) repeat families to at least the Order level, with 70 to the Superfamily level, corresponding to almost 6 million individual repeats summing to almost 1.5 Gb.
Bar charts showing the number of unknown repeats classified by Terrier for the northern krill at probability thresholds of 0.7 and 0.9. The stricter threshold of 0.9 results in fewer repeats being classified.
Figure 9: Extra classifications by Terrier at different probability threshold of previously unclassified repeat families from northern krill. The percentage of sequences remaining unknown is labeled for each threshold.
@@ 7 Conclusion
In this study, we introduce Terrier, a comprehensive software package designed for the precise classification of repeats from DNA sequences. Terrier distinguishes itself by incorporating a significantly greater number of prediction categories than other comparable methods, providing a finer level of detail for repeat classification. This expanded classification system enables more accurate and comprehensive assessments, especially when dealing with complex, diverse, or poorly characterized genomic regions.
@@
Terrier substantially reduced the number of unclassified repeats for experimental data of flatworms, amphibians and northern krill, demonstrating its effectiveness for large, highly repetitive genomes.
Compared to other state-of-the-art deep learning classifiers, Terrier demonstrates superior performance in terms of both classification accuracy and the breadth of sequences it can classify. The tool’s high precision, combined with its broad applicability across a range of species and repeat types, positions it as a leading tool for repeat classification. Terrier integrates seamlessly into existing repeat annotation workflows by running between the standard tools RepeatModeler and RepeatMasker. These capabilities make Terrier an essential tool for researchers studying repeat-driven evolution, genomic instability, and other areas where a detailed understanding of repeats is crucial. By advancing the field of repeat classification, Terrier enables more accurate genomic annotations and supports the exploration of complex genomic features in both model and non-model organisms.
@@ 8 Data availability
The software is available under an Apache 2.0 Open Source License and can be downloaded and installed from GitHub (https://github.com/rbturnbull/terrier) and from the Python Package Index (https://pypi.org/project/bio-terrier/). Weights for the model are available with the 0.2.0 release of the software and this is automatically downloaded when first using the model. We also provide a notebook which can be launched in Google Colab to run Terrier on a cloud GPU and to download the results.
The experimental amphibian and flatworm data underlying this article are available in FigShare, at https://doi.org/10.26188/28578992. New repeat classifications for northern krill are available in FigShare, at https://doi.org/10.26188/29226188.
@@ 10 Author contributions statement
R.T. conceived and wrote the Terrier software and performed the training. R.T., N.Y. and T.K. prepared the training dataset. R.T. and E.T. ran the software packages on the test datasets. N.Y. prepared the flatworm experimental data. T.K. and E.T. prepared the amphibian experimental data. All reviewed the manuscript.
11 Key Points
  • Terrier is a deep learning model trained on the extensive Repbase library, designed to improve repeat classification accuracy across a broad range of eukaryotic species.
  • Terrier outperforms existing tools such as DeepTE, TERL, and TEclass2 on test datasets from rice, fruit flies, humans, and mice.
  • We demonstrate its effectiveness on experimental data from non-model organisms of flatworms and amphibians. These datasets are available on FigShare under a Creative Commons open access license.
  • Terrier is available under the Apache 2.0 open source license, along with trained model weights. Documentation includes instructions to reproduce the results.
12 Funding
This project benefited from Australian Research Council grants FT190100462 and LP200301370 awarded to L.F.S.
@@ 13 Acknowledgments
This research was supported by The University of Melbourne’s Research Computing Services. We acknowledge the help of Priyanka Pillai, Swetha Gopikumar Sreeja and Rafsan Al Mamun.